OpenStudy (goformit100):

(a.) 8 men can finish a piece of work in 40 days. Then if 10 men work, then the work will be completed in how many days ? (b.) 8 pen costs 40 rupees. then 10 pen costs how many rupees ? ((At last compare between the procedures used in both the questions. Do they vary.)) HOW ? WHY ?

OpenStudy (goformit100):

Does it seems logically correct ? @dan815

OpenStudy (dan815):

ya no lol wait

OpenStudy (anonymous):

Yes in first case time is the cost of work that is it costs 40 days to complete work W by 8 men => 40 = W/8, so W= 320 man days, for 10 men it would take W/10 = 32 days In the second case the cost of 8 pens is 40 so cost per pen is 40/8 = 5, 10 pens cost 50

OpenStudy (dan815):

= 40*8/10

OpenStudy (goformit100):

Why using different procedures ?

OpenStudy (dan815):

and b whatever he said

OpenStudy (goformit100):

Dosent it seems wrong in mathematical point of view /

OpenStudy (dan815):

no why

OpenStudy (dan815):

oh umm i see the confusion yah if units are not considered

OpenStudy (anonymous):

lol he invited whole OS here

OpenStudy (john_es):

This is a problem of direct and indirect proportionality. \[y=kx\] for direct proportionality. \[y=\frac{k}{x}\] for indirect propotionality. (a) is indirect (the increase of some quantity provoke the decreasing of other quantity). (b) is direct, for bigger x bigger y.

OpenStudy (goformit100):

Don't we here use UNITARY METHOD ?

OpenStudy (somy):

is this question logical or mathematical? @goformit100

OpenStudy (somy):

8 men finish work in 40 days 10 man have extra 2 people so 2*40= 80/8= 10 days, so 40-10= 30 days

OpenStudy (somy):

just guessing lol

ganeshie8 (ganeshie8):

@John_ES 's reply is all you need @goformit100

OpenStudy (anonymous):

one is about time and other one about cost hw can they be same o_O

OpenStudy (anonymous):

increased number of men will reduce the time increased number of pens will increase the value

OpenStudy (anonymous):

what exactly is the confusion o_o

OpenStudy (somy):

in cost 10*40/8= 50 rup

OpenStudy (anonymous):

WE LOVE THE CROSS Multiplication BUT NOT THAT MUCH HEHE :P

OpenStudy (anonymous):

@aajugdar no confusion hehe

OpenStudy (anonymous):

hah :P

OpenStudy (mrnood):

@John_ES @goformit100 John_ES has the complete and exact answer - no other discussion is really required

OpenStudy (anonymous):

@goformit100 a) no time at all, the work has been finished already. b) 10 pens will cost 50 rupees.

OpenStudy (goformit100):

Okay, let me check the posts

OpenStudy (anonymous):

Hey how much % you secured in 12th

OpenStudy (goformit100):

I secured 66.2% in 12th @No.name

OpenStudy (anonymous):

Good , yesterday was the result right?

OpenStudy (goformit100):

No result was on 28th of May.

OpenStudy (anonymous):

You are from which board?

OpenStudy (goformit100):

I am from 'Central Board of Secondary Education'.

OpenStudy (anonymous):

CBSE

OpenStudy (goformit100):

Yes Sir.

OpenStudy (anonymous):

Are you good with chemistry , I am doing IIT

OpenStudy (anonymous):

@goformit100

OpenStudy (goformit100):

I am only better at Physical Chemistry than Inorganic and organic

OpenStudy (anonymous):

Okay , i am learning atomic structure i may have some doubts so shall i tag you

OpenStudy (goformit100):

Okay, I will surely give it a try.

OpenStudy (anonymous):

Okay , i will in the future

OpenStudy (goformit100):

But I personally have a video for you

OpenStudy (goformit100):

Can I ?

OpenStudy (anonymous):

Okay sir